the inverse of the function f(x)=10^x-10^-x/10^x+10^-x is

Grade:12th pass

1 Answers

Gowtham
6 years ago
let f(x)=y
solve the equation such that you will get the x as subject
after solving you will something like
x=(1/2)(log(1+y)/(1-y))
as f(x)=y → x=f-1(y)
f-1(y)=(1/2)(log(1+y)/(1-y))
so y can be replaced with x